Suppose that Wind Em Corp. currently has the balance sheet shown below, and that sales for the year just ended were $7.8 million. The firm also has a profit margin of 30 percent, a retention ratio of 15 percent, and expects sales of $8.8 million next year.

Assets Liabilities and Equity
Current assets $ 2,624,000 Current liabilities $ 3,622,320
Fixed assets 5,800,000 Long-term debt 1,900,000
Equity 2,901,680
Total assets $ 8,424,000 Total liabilities and equity $ 8,424,000

If all assets and current liabilities are expected to grow with sales, what amount of additional funds will Wind Em need from external sources to fund the expected growth? (Enter your answer in dollars not in millions.)  
Answer #1

Current years sales Next years sales Increase % in sales $7.8 $8.8 1 2.82% million million [($8.8m - $7.8m) / $8.8m] It isAlternatively, EFN can be calculated using formula: EFN = (S) - So) - (5. – S.)=(PM/S.XD) where • So = Current Sales, $7,800,